OpenStudy (anonymous):

the quadratic equation x2 - 4x + 57 = -5 by completing the square

OpenStudy (anonymous):

Move 57 to the other side of the eq so you have x^2-4x =-62

OpenStudy (anonymous):

Then take half of the "b" term, -4/2=-2. Square it, 4. Add to both sides of eq & factor X,^2-4x+4=-62

OpenStudy (anonymous):

(X-2)(X-2) or (X-2)^2=-62 Find the sq rt of each side. X-2=sq rt of -62 X-2 = i(sq rt 62) x = 2+i(Sq rt 62)